# Learning to Pursue AC Optimal Power Flow Solutions with Feasibility Guarantees

### v1

@@ II-C Proposed Mathematical Framework and Implementations
Figure 1: (Left) Feedback-based online implementation leveraging measurements from the network. (Center) Offline implementation with power-flow solver. (Right) Design process.
We design $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ based on CBF tools [32] and the safe gradient flow [23, 24]. In particular, $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is given by:
@@
$\arg\min_{\boldsymbol{z}\in\mathbb{R}^{2G}}\|\boldsymbol{z}+\nabla C_{p}(\boldsymbol{u})+\boldsymbol{J}_{v}(\boldsymbol{u};\boldsymbol{s}_{l})^{\top}\nabla C_{v}(\boldsymbol{\nu})\|^{2}$
[…]
@@
where $\boldsymbol{J}_{\ell_{i}}(p_{i},q_{i})$ is the Jacobian of $(p_{i},q_{i})\mapsto\ell_{i}(p_{i},q_{i})$, $\beta>0$ is a design parameter, and $\eta>0$ is the controller gain and is a design parameter. As discussed in [23], the controller in (6) serves as an approximation of the projected gradient flow. This approximation, which leverages CBF models, ensures that the feasible set of (II-B) is forward invariant. This invariance property is a key motivation behind initiating our design from (7), and it is supported by our recent work in [24], where the function $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta}_{F})$ was specifically designed to address an AC OPF problem with voltage constraints.
@@
$\triangle$ Model-based offline implementation: In this case, (5b) represents the solution to the AC power flow equations (2) via numerical methods. For example, given $\boldsymbol{s}_{l}$, voltages $\boldsymbol{v}$ can be found using the fixed-point method [28, 33]. This leads to the offline solution of (II-B) illustrated in Figure 1.
@@
Design and train a neural network $\mathcal{F}^{\textsf{NN}}$ to emulate $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ in (7) so that the algorithm (8): (a) converges to a solution $\boldsymbol{u}^{*}\in\textsf{U}^{\textsf{lm}}(\boldsymbol{\theta})$ of the AC OPF problem (II-B); (b) ensures that voltage, current, and DER constraints are satisfied at any time during the execution of the algorithm. $\Box$
The term “at any time” refers to the fact that the algorithm (8) is expected to produce points that are practically feasible for (II-B) even if it is terminated before convergence. This is a key for online AC OPF implementations as in Figure 1(left), and a desirable feature of offline methods. We provide some remarks to support our design approach.
II-D Motivations and Rationale
@@
  • For any given $\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta}$, $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is defined as the unique optimal solution of the convex QP (7). Moreover, under some mild assumptions, the mapping $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is locally Lipschitz (in all its arguments) [23, 24]. Therefore, in our approach, the neural network approximates a mapping that is continuous in its arguments. Accordingly, our learning problem is well posed.
@@
  • To approximate $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$, the inputs to the training are the voltages at the monitored nodes, the currents at the monitored lines, the current setpoints of the DERs, and the parameters $\boldsymbol{\theta}$. Here, the training does not include the loads $\boldsymbol{s}_{l}$ as inputs.
  • To approximate $\textsf{U}^{*}(\boldsymbol{s}_{l},\boldsymbol{\theta})$, $\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{l},\boldsymbol{\theta})$, or $\textsf{U}^{\textsf{kkt}}(\boldsymbol{s}_{l},\boldsymbol{\theta})$ directly, the inputs are the loads $\boldsymbol{s}_{l}$ throughout the network and the inputs $\boldsymbol{\theta}$. When the number of loads is larger than the controlled DERs and the monitored voltages and currents, this leads to more inputs in the training task.
@@ III-A Algorithmic Design
$F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta}):=\arg\min_{\boldsymbol{z}\in\mathbb{R}^{2G}}\|\boldsymbol{z}+\nabla C_{p}(\boldsymbol{u})+\boldsymbol{\Gamma}_{v}^{\top}\nabla C_{v}(\boldsymbol{\nu})\|^{2}$
[…]
@@
Similarly to (7),  (11) is a convex QP with a unique optimal solution. Moreover, from [35, Theorem 3.6], it follows that $\boldsymbol{u}\mapsto F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is locally Lipschitz over $\mathcal{B}(\boldsymbol{u},r_{1}):=\{\boldsymbol{z}:\|\boldsymbol{z}-\boldsymbol{u}\|<r_{1}\}$ of $\boldsymbol{u}$, for any $\boldsymbol{\xi}$ and $\boldsymbol{\theta}$.
@@
$\boldsymbol{y}$ $=\mathcal{F}^{\text{NN}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta}):=W^{(H)}\boldsymbol{\varphi}^{(H)}+b^{(H)},$ (12)
@@
For the training procedure, suppose that $N_{\text{train}}$ training points are available, and they are taken from a compact set $(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})\in\mathcal{C}_{\text{train}}\times\mathcal{E}_{\text{train}}\times\Theta_{\text{train}}$, where $\mathcal{C}_{\text{train}}$ is a superset of the feasible region of (II-B), $\mathcal{E}_{\text{train}}$ is an inflation of the set of operational voltages, and $\Theta_{\text{train}}$ is formed based on inverters’ operating conditions. Thus, each training point is given by the input $(\boldsymbol{u}^{(k)},\boldsymbol{\xi}^{(k)},\boldsymbol{\theta}^{(k)})$ and the corresponding output $\boldsymbol{y}^{(k)}=F_{\textsf{ln}}(\boldsymbol{u}^{(k)},\boldsymbol{v}^{(k)},\boldsymbol{\theta}^{(k)})$, for $k=1,\ldots,N_{\text{train}}$. Then, we consider minimizing the following loss function:
$\mathcal{L}(\boldsymbol{W},\boldsymbol{b}):=\frac{1}{N_{\text{train}}}\sum_{n=1}^{N_{\text{train}}}\left\|\boldsymbol{y}^{(k)}-\mathcal{F}^{\mathsf{NN}}(\boldsymbol{u}^{(k)},\boldsymbol{v}^{(k)},\boldsymbol{\theta}^{(k)})\right\|_{2}^{2}$ (13)
@@
We note that the training dataset can be generated offline by repeating step [S1] for a given set of values for voltages, currents, and DERs’ powers, or it can be formed online by collecting measurements from the distribution grid. The trained FNN $\mathcal{F}^{\mathsf{NN}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is then used in (8) to solve the AC OPF online (cf. Figure 1(left) or offline (cf. Figure 1(center)).
@@ III-B Online and Offline Implementations
On the other hand, the offline implementation of Figure 1(center) is tabulated as Algorithm 3. For this offline implementation, one needs information about the loads $\boldsymbol{s}_{l}$. In [S1a], solutions to the power flow (PF) equations can be identified using, for example, sweeping methods [26] or fixed-point methods [28]. In this implementation, the error $\boldsymbol{n}$ in (5a) and (8a) represents the numerical accuracy of the PF method. The algorithm is executed until convergence, or for a pre-scribed amount of time $t_{d}$.
@@ IV Experimental Results in a Distribution Feeder
We test the proposed method illustrated in Figure 1(left) – which in this section we refer to as neural network-based safe gradient flow (NN-SGF) in short - on a voltage regulation problem.
We consider the medium voltage network (20 kV) shown in Figure 2 (see [25]). The network contains photovoltaic (PV) inverters at selected buses capable of adjusting both active and reactive power. Each inverter $i\in\mathcal{G}$ injects power $\boldsymbol{u}_{i}=(p_{i},q_{i})\in\mathbb{R}^{2}$ within a feasible set:
@@
where $s_{n,i}$ represents the inverter’s nominal apparent power rating, randomly selected from the set {490, 620, 740} kVA to capture the range of deployment scales observed in practice. The upper bound $p_{\max,i}$ denotes the maximum available active power at time $t$. Together, the pair $\boldsymbol{\theta}_{u,i}=(p_{\max,i},s_{n,i})$ defines the parameterization of the set $\mathcal{C}_{i}(\boldsymbol{\theta}_{u,i})$. This limit is consistent with practical deployment settings and in accordance with IEEE Std 1547-2018. We also assume that $p_{\max,i}$ is known at the DER level (via maximum power point tracking). The cost function in the AC OPF is defined as $C_{p}(\boldsymbol{u})=\sum_{i\in\mathcal{G}}c_{p}\left(\frac{s_{n,i}-p_{i}}{s_{n,i}}\right)^{2}+c_{q}\left(\frac{q_{i}}{s_{n,i}}\right)^{2}$, with $c_{p}=3$ and $c_{q}=1$. This cost function aims to minimize active power curtailment and inverter losses. The first term promotes operation near the available active power, while the second penalizes reactive power usage, which contributes to higher current magnitudes and associated Joule losses.
@@
$\triangle$ (s1) A solution of the AC OPF every 10 seconds to match the granularity of the Open Power System Data. Here, we use the nonlinear branch flow model [37] and the solver IPOPT. We refer to this case as batch optimization (BO).
$\triangle$ (s2) Our solution strategy in (8) deployed in the online feedback-based configuration in Figure 1(left). Here, we run one iteration of the NN-SGF time a measurement is collected (as in standard feedback optimization methods).
@@ IV-A Dataset Generation and Training
[…]
@@
Learning Model, Training, and Evaluation. We train a fully connected FNN $\mathcal{F}^{\textsf{NN}}:\mathbb{R}^{2G}\times\mathbb{R}^{M+L}\times\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}^{2G}$; the FNN has an architecture of the form $[3G+M,h,h,h,2G]$, with hidden width set to $h=\alpha(3G+M)$; we use $\alpha=2$, yielding $h=524$ for $G=84$ and $M=10$. The network is implemented in PyTorch and trained offline using the Adam optimizer [^4] See: https://docs.pytorch.org/docs/stable/generated/torch.optim.Adam.html with learning rate 0.001, batch size 256, dropout 0.2, and up to 500 epochs with early stopping based on validation loss from a 10% held-out subset. The loss function is the mean squared error (MSE) $\mathcal{L}(\theta)=\frac{1}{|\mathcal{D}_{\text{train}}|}\sum_{n,k}\left\|\mathcal{F}^{\textsf{NN}}(\boldsymbol{u}^{(n,k)},\boldsymbol{\xi}^{(n,k)},\boldsymbol{\theta}^{(n)};\theta)-\boldsymbol{y}^{(n,k)}\right\|_{2}^{2}$ over the dataset $\mathcal{D}_{\text{train}}=\left\{(\boldsymbol{u}^{(n,k)},\boldsymbol{\xi}^{(n,k)},\boldsymbol{\theta}^{(n)},\boldsymbol{y}^{(n,k)})\right\}_{n,k}$. Performance on a disjoint test set $\mathcal{D}_{\text{test}}$ is evaluated using the MSE $\ell_{2}$-norm prediction error, defined as $\varepsilon^{\textsf{NN}}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{n,k}\left\|\mathcal{F}^{\textsf{NN}}(\boldsymbol{u}^{(n,k)},\boldsymbol{\xi}^{(n,k)},\boldsymbol{\theta}^{(n)})-\boldsymbol{y}^{(n,k)}\right\|_{2}^{2}$ ,yielding $\varepsilon^{\textsf{NN}}=1.7\times 10^{-6}$ and $\mathrm{RMSE}=0.0013$. During online deployment, the FNN runs in inference mode at each control step $t$, with sampling interval $\Delta t=10$ seconds (to match the variability of the load and PV data). Given current DER setpoints $\boldsymbol{u}(t)$, measurements $\boldsymbol{\xi}(t)$, and parameters $\boldsymbol{\theta}(t)$, the update is approximated as $\boldsymbol{u}(t+\Delta t)=\boldsymbol{u}(t)+\eta\Delta t\cdot\mathcal{F}^{\textsf{NN}}(\boldsymbol{u}(t),\boldsymbol{\xi}(t),\boldsymbol{\theta}(t))$, with $\eta=0.02$; setpoints are restricted to the set $\mathcal{C}$, via a projection, if not feasible to reflect hardware constraints. Given the setpoints, updated voltage magnitudes are computed via AC power flow using pandapower, yielding the new vector $\boldsymbol{\xi}(t+\Delta t)$ for the next control step.

### v2

@@ II-B Proposed Mathematical Framework and Implementations
Figure 1: (Left) Feedback-based online implementation, leveraging measurements of voltages and currents from the power distribution network. (Center) Offline implementation, where in the iterative process voltages and currents are computed using a power-flow solver. In both cases, $\xi$ collects the voltages and line-currents $(|\boldsymbol{v}|,|\boldsymbol{i}|)$. (Right) Design process, where the linear approximation of the power flow equations is utilized to lower the computational complexity of the training.
We design $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ based on the safe gradient flow (SGF) [1, 11]. In particular, $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is given by:
@@
$\arg\min_{\boldsymbol{z}\in\mathbb{R}^{2G}}\|\boldsymbol{z}+\nabla\phi_{p}(\boldsymbol{u})+\boldsymbol{J}_{v}(\boldsymbol{u};\boldsymbol{s}_{l})^{\top}\nabla\phi_{v}(\boldsymbol{\nu})\|^{2}$
$\hskip 28.45274pt\textrm{s.t.}-\boldsymbol{J}_{v}(\boldsymbol{u};\boldsymbol{s}_{l})^{\top}\boldsymbol{z}\leq-\beta\left(\mathbf{1}\underline{V}-\boldsymbol{\nu}\right)$ (7)
@@
where $\boldsymbol{J}_{\ell_{i}}(p_{i},q_{i})$ is the Jacobian of $(p_{i},q_{i})\mapsto\ell_{i}(p_{i},q_{i})$, $\beta>0$ is a design parameter, and $\eta>0$ is the controller gain and is a design parameter (inequalities are taken entry-wise). As discussed in [1], the controller in (6) serves as an approximation of the projected gradient flow. This approximation, which leverages control barrier function (CBF) tools, ensures that the feasible set of (II-A) is forward invariant. This invariance property is a key motivation behind initiating our design from (7), and it is supported by our recent work in [11], where the function $F(\boldsymbol{u},\boldsymbol{\xi},{\boldsymbol{\theta}})$ was specifically designed to address an AC OPF problem with voltage constraints.
@@
$\triangle$ Model-based offline implementation: In this case, (5b) represents the solution to the AC power flow equations (2) via numerical methods. For example, given $\boldsymbol{s}_{l}$, voltages $\boldsymbol{v}$ can be found using the fixed-point method [6, 10]. This leads to the offline solution of (II-A) illustrated in Figure 1.
@@
Train a neural network $\mathcal{F}^{\textsf{NN}}$ to emulate $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ in (7) so that the algorithm (8): (a) converges to a solution $\boldsymbol{u}^{*}\in\textsf{U}^{\textsf{lm}}(\boldsymbol{\theta})$ of the AC OPF problem (II-A); (b) ensures that voltage, current, and DER constraints are satisfied at any time during the execution of the algorithm. $\Box$
The term “at any time” refers to the fact that the algorithm (8) is expected to produce points that are practically feasible for (II-A) even if it is terminated before convergence. This is a key for online AC OPF implementations as in Figure 1(left), and a desirable feature of offline methods. We provide some remarks to support our design approach.
II-C Motivations and Rationale
@@
  • For any given $\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta}$, $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is defined as the unique optimal solution of the convex QP (7). Moreover, under some mild assumptions, the mapping $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is locally Lipschitz (in all its arguments) [1, 11]. Therefore, in our approach, the neural network approximates a mapping that is continuous in its arguments.
@@
  • To approximate $\mathcal{F}^{\textsf{NN}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$, the inputs to the training are: (i) the monitored voltage and/or line-current magnitudes collected in $\boldsymbol{\xi}$, (ii) the current DER setpoints $\boldsymbol{u}$, and (iii) the parameters $\boldsymbol{\theta}$ (see Fig. 1). We recall that $\boldsymbol{\xi}$ is obtained from measurements in the feedback-based online implementation or from numerical power-flow solutions in the offline implementation. The loads $\boldsymbol{s}_{\ell}$ are not included in the feature vector used for training.
  • To approximate $\textsf{U}^{*}(\boldsymbol{s}_{\ell},\boldsymbol{\theta})$, $\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{\ell},\boldsymbol{\theta})$, or $\textsf{U}^{\textsf{kkt}}(\boldsymbol{s}_{\ell},\boldsymbol{\theta})$ directly, the inputs are the system-wide loads $\boldsymbol{s}_{\ell}$. Thus, when the number of load buses is large (e.g., larger than the number of controlled DERs and the number of monitored voltage/current measurements), the dimension of the corresponding input vector grows with $\dim(\boldsymbol{s}_{\ell})$, and can be substantially larger than the input dimension required to learn $F(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$, which depends on the current setpoints $\boldsymbol{u}$ and on $\boldsymbol{\xi}=\big(|\boldsymbol{v}|,|\boldsymbol{i}|\big)$.
@@ III-A Algorithmic Design
$F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta}):=\arg\min_{\boldsymbol{z}\in\mathbb{R}^{2G}}\|\boldsymbol{z}+\nabla\phi_{p}(\boldsymbol{u})+\boldsymbol{\Gamma}_{v}^{\top}\nabla\phi_{v}(\boldsymbol{\nu})\|^{2}$
$\hskip 85.35826pt\textrm{s.t.}-\boldsymbol{\Gamma}_{v}^{\top}\boldsymbol{z}\leq-\beta\left(\mathbf{1}\underline{V}-\boldsymbol{\nu}\right)$ (11)
@@
Similarly to (7),  (11) is a convex QP with a unique optimal solution. Moreover, from [26, Theorem 3.6], it follows that $\boldsymbol{u}\mapsto F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is locally Lipschitz over $\mathcal{B}(\boldsymbol{u},r_{1}):=\{\boldsymbol{z}:\|\boldsymbol{z}-\boldsymbol{u}\|<r_{1}\}$ of $\boldsymbol{u}$, for any $\boldsymbol{\xi}$ and $\boldsymbol{\theta}$.
@@
$\boldsymbol{y}$ $=\mathcal{F}^{\textsf{NN}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta}):=W^{(H)}\boldsymbol{\varphi}^{(H)}+b^{(H)},$ (12)
@@
For the training procedure, suppose that $N_{\text{train}}$ training points are available, and they are taken from a compact set $(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})\in\mathcal{C}_{\text{train}}\times\mathcal{E}_{\text{train}}\times\Theta_{\text{train}}$, where $\mathcal{C}_{\text{train}}$ is a superset of the feasible region of (II-A), $\mathcal{E}_{\text{train}}$ is an inflation of the set of operational voltages, and $\Theta_{\text{train}}$ is formed based on inverters’ operating conditions. Thus, each training point is given by the input $(\boldsymbol{u}^{(k)},\boldsymbol{\xi}^{(k)},\boldsymbol{\theta}^{(k)})$ and the corresponding output $\boldsymbol{y}^{(k)}=F_{\textsf{ln}}(\boldsymbol{u}^{(k)},\boldsymbol{\xi}^{(k)},\boldsymbol{\theta}^{(k)})$, for $k=1,\ldots,N_{\text{train}}$. Then, we consider minimizing the following loss function:
$\hskip-5.69046pt\mathcal{L}(\boldsymbol{W},\boldsymbol{b}):=\frac{1}{N_{\text{train}}}\sum_{n=1}^{N_{\text{train}}}\left\|\boldsymbol{y}^{(k)}-\mathcal{F}^{\mathsf{NN}}(\boldsymbol{u}^{(k)},\boldsymbol{\xi}^{(k)},\boldsymbol{\theta}^{(k)})\right\|_{2}^{2}$ (13)
@@
We note that the training dataset can be generated offline by repeating the data-generation loop in Algorithm 1 (steps 1–5) for a given set of values for voltages, currents, and DERs’ powers, or it can be formed online by collecting measurements from the distribution grid. The trained FNN $\mathcal{F}^{\mathsf{NN}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ is then used in (8) to solve the AC OPF online (cf. Figure 1(left) or offline (cf. Figure 1(center)).
@@ III-B Online and Offline Implementations
On the other hand, the offline implementation of Figure 1(center) is tabulated as Algorithm 3. For this offline implementation, one needs information about the loads $\boldsymbol{s}_{l}$. In Algorithm 3 (step 5), solutions to the power flow (PF) equations can be identified using, for example, sweeping methods [24] or fixed-point methods [6]. In this implementation, the error $\boldsymbol{n}$ in (5a) and (8a) represents the numerical accuracy of the PF method. The algorithm is executed until convergence, or for a pre-scribed amount of time $t_{d}$.
@@ IV Experimental Results in a Distribution Feeder
We test the proposed method illustrated in Figure 1(left) – which in this section we refer to as neural network-based safe gradient flow (NN-SGF) in short - on a voltage regulation problem. We consider the medium voltage network (20 kV) shown in Figure 2 (see [35]). We define all per-unit quantities using $V_{\text{base}}=20~\text{kV}$ and $S_{\text{base}}=10~\text{MVA}$, yielding $Z_{\text{base}}=V_{\text{base}}^{2}/S_{\text{base}}=40~\Omega$. The network contains photovoltaic (PV) inverters at selected buses capable of adjusting both active and reactive power. Each inverter $i\in\mathcal{G}$ injects power $\boldsymbol{u}_{i}=(p_{i},q_{i})\in\mathbb{R}^{2}$ within a feasible set:
@@
where $s_{n,i}$ represents the inverter’s nominal apparent power rating, randomly selected from the set {490, 620, 740} kVA (i.e., $s_{n,i}/S_{\text{base}}\in\{0.049,\,0.062,\,0.074\}$ p.u.) to capture the range of deployment scales observed in practice. The upper bound $p_{\max,i}$ denotes the maximum available active power at time $t$. Together, the pair $\boldsymbol{\theta}_{u,i}=(p_{\max,i},s_{n,i})$ defines the parameterization of the set $\mathcal{C}_{i}(\boldsymbol{\theta}_{u,i})$. This limit is consistent with practical deployment settings and in accordance with IEEE Std 1547-2018. We also assume that $p_{\max,i}$ is known at the DER level (via maximum power point tracking). The cost function in the AC OPF is defined as $\phi_{p}(\boldsymbol{u})=\sum_{i\in\mathcal{G}}c_{p}\left(\frac{s_{n,i}-p_{i}}{s_{n,i}}\right)^{2}+c_{q}\left(\frac{q_{i}}{s_{n,i}}\right)^{2}$, with $c_{p}=3$ and $c_{q}=1$. Here, the first term aims to minimize active power curtailment and inverter losses, whereas the latter discourages reliance on reactive power, as it leads to increased current levels and energy losses.
@@
$\triangle$ (s1) A solution of the AC OPF every 10 seconds to match the granularity of the Open Power System Data. Here, we use the nonlinear branch flow model [4] and the solver IPOPT. We refer to this case as batch optimization (BO), meaning that at every time we solve the nonlinear AC OPF problem to convergence.
$\triangle$ (s2) Our solution strategy in (8) is deployed in the online feedback-based configuration in Figure 1 (left). Here, we run one iteration of NN–SGF each time a measurement is collected (as in standard feedback optimization methods). Here, “one iteration” means a single forward evaluation of $\mathcal{F}^{\textsf{NN}}\!\big(\boldsymbol{u}(t_{k}),\boldsymbol{\xi}(t_{k}),\boldsymbol{\theta}(t_{k})\big)$ using the latest measurement at time $t_{k}$; the setpoints are then held until the next measurement arrives.
@@ IV-A Dataset Generation and Training
We simulate $N_{\text{training}}=6{,}000$ independent operating conditions, each defined by a distinct grid configuration with varying load profiles, DER capacities, and voltage regulation constraints. Each condition is associated with a randomly sampled time instant $t_{n}\sim\mathcal{U}(t_{\min},t_{\max})$, where $t_{\min}=\text{06{:}00}$, $t_{\max}=\text{20{:}00}$. For each sampled time $t_{n}$, we simulate power flow using pandapower to obtain voltages $\boldsymbol{\nu}(t_{n})=\{V_{j}(t_{n})\}_{j\in\mathcal{M}}$, along with DER setpoints $\{p_{i}(t_{n}),q_{i}(t_{n})\}_{i\in\mathcal{G}}$, constraint parameters $\{p_{\max,i}(t_{n}),s_{n,i}\}_{i\in\mathcal{G}}$, and voltage bounds $\{\underline{V}_{j},\overline{V}_{j}\}_{j\in\mathcal{M}}$. For each operating condition, we run $N_{\text{iter}}=10$ iterations of $F_{\textsf{ln}}(\boldsymbol{u},\boldsymbol{\xi},\boldsymbol{\theta})$ using a forward Euler discretization, with $\eta=0.2$. At each iteration $k\in\{1,\ldots,N_{\text{iter}}\}$, we record the DER setpoints $\boldsymbol{u}^{(n,k)}=\{p_{i}^{(k)}(t_{n}),q_{i}^{(k)}(t_{n})\}_{i\in\mathcal{G}}$, the voltage measurements $\boldsymbol{\nu}^{(n,k)}=\{V_{j}^{(k)}(t_{n})\}_{j\in\mathcal{M}}$, and the constraint parameters $p_{\max,i}(t_{n})$, $s_{n,i}$, $\underline{V}_{j}$, and $\overline{V}_{j}$. We also extract the current magnitudes $\boldsymbol{\iota}^{(n,k)}\in\mathbb{R}^{L}$ at monitored lines. These are used to construct the state vector $\boldsymbol{\xi}^{(n,k)}=[(\boldsymbol{\nu}^{(n,k)})^{\top},\,(\boldsymbol{\iota}^{(n,k)})^{\top}]^{\top}\in\mathbb{R}^{M+L}$, consistent with the online and offline implementation setup. The full constraint parameter vector is denoted $\boldsymbol{\theta}^{(n)}=(\boldsymbol{\theta}_{u,1}^{(n)},\ldots,\boldsymbol{\theta}_{u,G}^{(n)},\underline{V},\overline{V},\overline{I})$, and is treated as fixed across SGF iterations at time $t_{n}$. The control update is computed as $\boldsymbol{y}^{(n,k)}=F_{\textsf{ln}}(\boldsymbol{u}^{(n,k)},\boldsymbol{\xi}^{(n,k)},\boldsymbol{\theta}^{(n)})\in\mathbb{R}^{2G}$ and used as the training label. The corresponding input vector $\mathbf{x}^{(n,k)}\in\mathbb{R}^{3G+M}$ is constructed by concatenating the normalized active power deviations $\left\{(p_{i}^{(k)}(t_{n})-p_{\max,i}(t_{n}))/s_{n,i},i\in\mathcal{G}\right\}$, normalized reactive powers $\left\{q_{i}^{(k)}(t_{n})/s_{n,i},i\in\mathcal{G}\right\}$, normalized voltage magnitudes $\left\{(V_{j}^{(k)}(t_{n})-\underline{V}_{j})/(\overline{V}_{j}-\underline{V}_{j}),j\in\mathcal{M}\right\}$, and the active DER limits $\left\{p_{\max,i}(t_{n}),i\in\mathcal{G}\right\}$. This normalization ensures all features lie in comparable ranges and are scaled relative to their physical limits (e.g., $s_{n,i}$ and voltage bounds), improving numerical stability. This process results in a dataset of $N_{\text{train}}=N_{\text{training}}\cdot N_{\text{iter}}=60{,}000$ input-output pairs $\mathcal{D}_{\text{train}}=\left\{\left(\mathbf{x}^{(n,k)},\mathbf{y}^{(n,k)}\right)\right\}_{n,k}\subset\mathbb{R}^{3G+M}\times\mathbb{R}^{2G}$. For evaluation, we construct a disjoint test set $\mathcal{D}_{\text{test}}=\left\{\left(\mathbf{x}^{(n,k)},\mathbf{y}^{(n,k)}\right)\right\}_{n,k}\subset\mathbb{R}^{3G+M}\times\mathbb{R}^{2G}$ using $N_{\text{testing}}=1000$ randomly sampled times $t_{n}\sim\mathcal{U}(\text{06{:}00},\;\text{20{:}00})$, with the same SGF update procedure repeated for each test case. We ensure that $\mathcal{D}_{\text{train}}\cap\mathcal{D}_{\text{test}}=\emptyset$.
@@
Learning Model, Training, and Evaluation. We train a fully connected FNN $\mathcal{F}^{\textsf{NN}}:\mathbb{R}^{2G}\times\mathbb{R}^{M+L}\times\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}^{2G}$; the FNN has an architecture of the form $[3G+M,h,h,h,2G]$, with hidden width set to $h=\alpha(3G+M)$; here, $\alpha$ scales the hidden width and thus directly sets the model capacity. We use $\alpha=2$, yielding $h=524$ for $G=84$ and $M=10$. The network is implemented in PyTorch and trained offline using the Adam optimizer [^4] See: https://docs.pytorch.org/docs/stable/generated/torch.optim.Adam.html with learning rate 0.001, batch size 256, dropout 0.2, and up to 500 epochs with early stopping based on validation loss from a 10% held-out subset. All experiments were run on a laptop with a 16-core Intel Core Ultra 7 255H CPU and 32 GB of RAM. Under this setup, training NN–SGF takes approximately 75 minutes. The loss function is the mean squared error (MSE) $\mathcal{L}(\theta)=\frac{1}{|\mathcal{D}_{\text{train}}|}\sum_{n,k}\left\|\mathcal{F}^{\textsf{NN}}(\boldsymbol{u}^{(n,k)},\boldsymbol{\xi}^{(n,k)},\boldsymbol{\theta}^{(n)};\theta)-\boldsymbol{y}^{(n,k)}\right\|_{2}^{2}$ over the dataset $\mathcal{D}_{\text{train}}=\left\{(\boldsymbol{u}^{(n,k)},\boldsymbol{\xi}^{(n,k)},\boldsymbol{\theta}^{(n)},\boldsymbol{y}^{(n,k)})\right\}_{n,k}$. Performance on a disjoint test set $\mathcal{D}_{\text{test}}$ is evaluated using the MSE $\ell_{2}$-norm prediction error, defined as $\varepsilon^{\textsf{NN}}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{n,k}\left\|\mathcal{F}^{\textsf{NN}}(\boldsymbol{u}^{(n,k)},\boldsymbol{\xi}^{(n,k)},\boldsymbol{\theta}^{(n)})-\boldsymbol{y}^{(n,k)}\right\|_{2}^{2}$ ,yielding $\varepsilon^{\textsf{NN}}=1.7\times 10^{-6}$ and $\mathrm{RMSE}=0.0013$. During online deployment, the FNN runs in inference mode at each control step $t$, with sampling interval $\Delta t=10$ seconds (to match the variability of the load and PV data). Given current DER setpoints $\boldsymbol{u}(t)$, measurements $\boldsymbol{\xi}(t)$, and parameters $\boldsymbol{\theta}(t)$, the update is approximated as $\boldsymbol{u}(t+\Delta t)=\boldsymbol{u}(t)+\eta\Delta t\cdot\mathcal{F}^{\textsf{NN}}(\boldsymbol{u}(t),\boldsymbol{\xi}(t),\boldsymbol{\theta}(t))$, where $\eta$ is the controller gain of the safe gradient flow controller (higher values lead to more aggressive updates), and it is set to $\eta=0.02$; setpoints are restricted to the set $\mathcal{C}$, via a projection, if not feasible to reflect hardware constraints. Given the setpoints, updated voltage magnitudes are computed via AC power flow using pandapower, yielding the new vector $\boldsymbol{\xi}(t+\Delta t)$ for the next control step.